INFRARED SPECTROSCOPY OF THREE ADENINE-MAGNESIUM ISOMERS SOLVATED IN SUPER-FLUID HELIUM DROPLETS

\underline{P. L. STILES}, M. Y. Choi, and R. E. MILLER, Department of Chemistry, The University of North Carolina at Chapel Hill, Chapel Hill, NC 27599.

Adenine, a DNA base, has been complexed with a single magnesium atom within superfluid helium droplets. The N9H stretch, symmetric and asymmetric NH_{2} stretching modes have been measured and compared to the adenine monomer. It was found that three separate isomers exist in the helium droplets environment, specifically, magnesium binds to the N1, N3 and N7 sites on the adenine. Vibrational bands are unambiguously characterized using vibrational transition moment angle (VTMA) analysis.